OpenStudy (anonymous):

Verify the given linear approximation at a = 0. Then determine the values of x for which the linear approximation is accurate to within 0.1. (Enter your answer using interval notation. Round your answer to two decimal places.) tan x ≈ x

OpenStudy (freckles):

So you want the magnitude of the difference of tan(x) and x to be less than or equal to 0.1 .

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

like the closest one to 0.1 but no 0.1. like 0.09

OpenStudy (freckles):

\[|\tan(x)-x| \le 0.1 \\ -0.1 \le \tan(x)-x \le 0.1 \\ x-0.1 \le \tan(x) \le x+0.1 \] so you are saying within doesn't include .1 ?

OpenStudy (anonymous):

yes, it is not include it.

OpenStudy (anonymous):

I did that same thing but like i dont know what to do then

OpenStudy (freckles):

\[|\tan(x)-x| < 0.1 \\ -0.1 < \tan(x)-x < 0.1 \\ x-0.1 <\tan(x) < x+0.1\]

OpenStudy (anonymous):

because a=0 so i tried [-0.1 , 0.1] and it told me it doesn't include 0.1

OpenStudy (freckles):

yeah I don't know either at this second

OpenStudy (anonymous):

Do you have chegg?

OpenStudy (freckles):

i don't know what chegg is but I know we should at least look in the interval (-pi/2,pi/2)

OpenStudy (freckles):

since at x=-pi/2 and x=pi/2 for y=tan(x) we have vertical asymptotes...

OpenStudy (freckles):

\[\frac{-\pi}{2} < \tan(x)<\frac{\pi}{2}\]

OpenStudy (anonymous):

yeah i tried that. its incorrect

OpenStudy (freckles):

Well I'm not done playing with this idea I

OpenStudy (anonymous):

i am here because i am desesperate haha i have tried everything i know

OpenStudy (aum):

If you are allowed to use a graphing calculator, click on the link and then click on the graph to enlarge it: http://awesomescreenshot.com/09d3p1fv1b

OpenStudy (anonymous):

I am not supposed to.. but i mean, everyone tells me its the only way! Thanks, i will try that and let you know

OpenStudy (aum):

Rounded to two decimal places: (-0.63, 0.63)

OpenStudy (anonymous):

It is correct! Thank you so much, you just literally saved my life haha. BTW, what is that website with that graphic calc. I want it

OpenStudy (aum):

www.desmos.com

OpenStudy (aum):

Glad that was accepted. You are welcome.